I have a homework assignment, and I'm supposed to "factor the algebraic expressions completely." I've been able to figure out everything but one.
Here's the one giving me trouble:
x^8 - 1
I have no idea where to begin! Any help would be greatly appreciated.

Answer by stanbon(75887)

x^8 - 1
The problem is in the form of a "difference of squares".
=(x^4-1)(x^4+1)
=(x^2-1)(x^2+1)(x^4+1)
=(x-1)(x+1)(x^2+1)(x^4+1)
==============
Cheers,
Stan H.